# NeSyGeo: A Neuro-Symbolic Framework for Multimodal Geometric Reasoning Data Generation

### v1

@@ 3.2 Symbolic Sequence Generation
To generate a Geo-DSL sequence, we introduce a step-action augmenter that iteratively synthesizes a sequence of statements, as detailed in Algorithm 1. Based on dataset preferences, we first configure the step count $N$, weight matrices $I$ and $A$ for selecting elements and actions with respective probabilities, and ranges $[l_{\text{min}},l_{\text{max}}]$ for lengths and $[\theta_{\text{min}},\theta_{\text{max}}]$ for angles. Then, the augmenter iteratively generates symbolic statements over $N$ steps. For each step, we randomly sample parameters $x,y,z$ and $\alpha$, select an element $v_{j}\in f_{v}$ using weights from $I$, and choose an action $a_{k}$ based on weights from $A$ (see Appendix I for action details). The new statement $s_{\text{new}}$ is then incorporated into the sequence $f_{s}$. Leveraging Geo-DSL’s symbolic definitions and well-defined actions, this approach ensures the validity and accuracy of each statement. Meanwhile, randomized elements, diverse action selections, and customizable hyperparameter preferences promote diversity.
1:Input: Step count $N$, Action weight matrix $A$, Element selection weight matrix $I$, Line length range $[l_{min},l_{max}]$, Angle range $[\theta_{min},\theta_{max}]$$\triangleright$ Set customizable hyperparameter
2:Output: Generated Geo-DSL statement sequence $f_{s}$.
5:for $i=1$ to $N$ do
6:     Randomly sample $x,y,z$ from $[l_{min},l_{max}]$.
7:     Randomly sample $\alpha$ from $[\theta_{min},\theta_{max}]$.
8:     Element $v_{j}$=Selected_elements($f_{v}$,$I$). $\triangleright$ Select element $v_{j}$ from $f_{v}$ randomly.
9:     Action $a_{k}$=Selected_action($v_{j}$,$A$). $\triangleright$ Select action $a_{k}$ based on the type of $v_{j}$ randomly
11:     $f_{s}$=Update($f_{s}$,$s_{new}$) $\triangleright$ Add the new statement to the sequence
12:     $f_{v}$=Update($f_{v}$,$s_{new}$). $\triangleright$ Update state space elements
13:end for
14:return $f_{s}$.
@@ 3.3 Informalization
Following the generation of sequences within the formal symbolic space, it is necessary to map them back to the text and image spaces for further processing and visualization. Our approach generates high-quality images and rigorous natural language, allocating information between them to compel MLLMs to leverage visual data effectively.
Figure 4: Reverse search and forward validation with expert LLMs
Visual image. For the visual space, our visualization engine parses each Geo-DSL statement as a geometric element to generate high-quality images using Matplotlib. The rigorous symbolic language space ensures the determinacy and uniqueness of the conversion process, thereby enabling precise image generation. Additionally, the engine produces images with detailed annotations absent from the text, requiring models to leverage images during problem-solving, enhancing their visual perception and ability to extract image-based information.
Natural language. We adopt a template-based transformation approach, parsing and mapping each symbolic statement to multiple predefined natural language templates. This produces a text-full version containing all details and captions of the image and a text-lite version retaining only essential conditions, not included in the image annotations, as the final form in our datasets. The text-lite version avoids intermodal information redundancy, while diverse templates ensure the validity and diversity of the synthesized condition text.
3.4 CoT Generation
To generate Q&A pairs with CoT reasoning, we utilize the strong reasoning abilities of expert LLMs to ensure diverse and reliable output. Using the text-full version generated as described in Section 3.3 as input, we develop a two-step process comprising reverse search and forward validation as shown in Figure 4. Prompts are detailed in Appendix G.
Reverse search. We use DeepSeek R1 [6] for reverse search, starting from the given conditions, iteratively exploring and deriving conclusions step by step, and ultimately producing the final conclusions at the end of the reasoning chain along with their corresponding answers. This reverse approach starts from known information and recursively builds a reasoning chain, reducing Q&A generation complexity and hallucinations. R1’s strong reasoning and exploration capabilities yield diverse conclusions, enriching the variety of Q&A pairs.
Forward validation. To ensure the correctness of Q&A pairs and generate step-by-step CoT reasoning, we re-input the questions and text-full conditions into DeepSeek V3, requesting both the reasoning process with final answer, and cross-validate these answers against those from R1 to include only consistent pairs in our final dataset. This process guarantees answer correctness and yields diverse, valid CoT without an extensive search space.
@@ 4 Experiments
We present a series of experiments to investigate the following four research questions:
Efficacy – To what extent does training on our synthesized dataset in both RL and SFT improve the geometric reasoning performances of several MLLMs?
Efficiency – Is the data generated by NeSyGeo more effective in yielding models with better performance compared to using data from existing automatic synthesis frameworks?
Diversity – Does the NeSyGeo framework effectively ensure diversity across both the text and image spaces of the generated synthetic geometric dataset?
Visual Effectiveness – Can our datasets compel models to effectively utilize visual information for enhanced understanding by appropriately distributing information between modalities?
@@ 4.1 Experimental Setup
Dataset. To synthesize a diverse dataset, we set hyperparameters for generating images by configuring the step count, length range, and angle range. The step count $N$ ranges from one to four. The line length is defined within the basic range $[l_{\text{min}},l_{\text{max}}]=[1,5]$, scalable by any multiple of 2 (e.g., $[4,20]$). The angle is constrained to multiples of 15° within $[15^{\circ},165^{\circ}]$, with increased weights assigned to special angles. We generate various types of weight matrices $A$ and $I$ by adjusting their corresponding values. This process yields a NeSyGeo-CoT dataset with 30k Q&A pairs and a NeSyGeo-Caption dataset with 70k Q&A pairs. Additional dataset statistics are provided in Appendix B.
Evaluation. Our evaluation is conducted on several benchmarks: the Test set of GeoQA[4], the Test_MINI set of MathVision[26], and the MathVerse[36]. For the MathVerse benchmark, we select the Vision Only, Vision Dominant, and Vision Intensive sets to better assess the visual perception and logical reasoning capabilities of MLLMs. We extract in-domain metrics from other datasets, including angle, area, length, and Plane Geometry, to effectively evaluate the models’ capabilities in geometric reasoning problems, in addition to the GeoQA dataset, which focuses entirely on plane geometry. For GeoQA, we employed hard-coded extraction for comparison, while other evaluations are assessed using the automated VLMEvalKit framework[8]. Appendix C provides additional experimental details and evaluation results.
@@ 4.2 Empirical Results
Efficacy: Training multiple MLLMs with NeSyGeo via both SFT and RL significantly enhances geometric problem-solving performances.
We first sample 4k samples from our NeSyGeo-CoT dataset and apply the Group Relative Policy Optimization (GRPO) algorithm to train two epochs with Deepseek R1’s format and answer rewards. The training code framework is based on VLM-R1 [22]. As shown in Tables 2, models achieve the best performance among the baselines after training. InternVL2.5-4B significantly improved in the angle knowledge domain with gains of 8.4 (MathVerse), 7.3 (GeoQA), and 5.3 (MathVision). Qwen2.5-VL-3B achieved a +15.8 performance boost in the area domain of MathVision. Notably, across all evaluated metrics, the InternVL2.5-4B model trained on the NeSyGeo dataset achieves performance on par with or superior to its 8B counterpart.
We also conducted SFT experiments, initially training on our NeSyGeo-Caption dataset to enhance the models’ perception of geometric images, followed by training on the NeSyGeo-CoT dataset to improve reasoning capabilities. The experiments were conducted on LLaMA-Factory [40] framework. Evaluation results on MathVerse (Vision Intensive) and GeoQA are presented in Table 3. The trained model demonstrates performance improvements over the base model on most metrics.
Efficiency: Under the same data budget, the generated data from our framework is better than that from popular automatic generation frameworks.
We randomly sampled 4k samples from MAVIS [37] and R-CoT [7], which are automatic frameworks in geometry problem generation. To ensure a fair comparison, we maintained consistent settings.
@@
Model Angle Area Length Angle Area Length Plane Geometry
@@
GeoQA Vision Intensive
Model Angle Area Length Plane Geometry
@@
Qwen2.5-VL-7B+NeSyGeo 71.8 (+2.4) 46.1 (+3.1) 23.1 (-4.4) 49.5 (+3.3) 46.7 (+2.6)
LLaVA-Next-7B 22.6 28.5 6.6 16.5 20.4
LLaVA-Next-7B+NeSyGeo 26.1 (+3.5) 30.6 (+2.1) 7.7 (+1.1) 19.2 (+2.7) 22.9 (+2.5)
As illustrated in Figure 5 and Table 2, while all datasets help the model improve over the baseline, our dataset outperformed others in most metrics, validating its superior performance.
Diversity: The dataset synthesized by our NeSyGeo framework exhibits high diversity in both text and visual features.
A critical challenge for automatic data synthesis methods is whether the dataset is sufficiently diverse to avoid quality degradation due to potential overfitting risks from inherent domain constraints. We employ t-SNE [24] dimensionality reduction for mapping in text space to evaluate the diversity across different methods. This analysis allows us to assess the diversity of the textual descriptions themselves. Given that in geometric problems, the text conditions depict specific visual elements, and the diversity observed in the text space also serves as a valuable indicator of the diversity in the corresponding visual diagrams. To ensure a fair comparison, we remove all prompts related to guiding large models, retaining only condition and question texts, and randomly sample 5k texts from each dataset. The results are illustrated in Figure 6. Our method and G-LLaVA [9] exhibit uniformly distributed features in the space, indicating low data overlap and high diversity. In contrast, R-CoT and MAVIS display varying degrees of clustered distribution, indicating more feature-similar samples. To directly assess the diversity of the visual features, we also performed t-SNE on image features extracted by ResNet, with detailed experimental results presented in Appendix C.
@@
Visual Perception: Models trained on NeSyGeo data shows modest gains over information-redundant datasets when textual shortcuts exist, yet achieves substantial improvements when image understanding is necessary. This indicates that ours enhances not only logical reasoning but also image perception and utilization of the model.
@@ 5 Conclusion
This paper introduces NeSyGeo, a neurosymbolic framework for automatically synthesizing multimodal geometric datasets. Our approach transforms the generation process into a controllable symbolic space using Geo-DSL, maps the symbolic representation back to image and natural language spaces via a conversion engine, and then utilizes LLMs for backwards search and forward solving to produce Q&A pairs. Using this framework, we construct the NeSyGeo-CoT and NeSyGeo-Caption datasets, totalling 100k samples. We also propose NeSyGeo-Test, a comprehensive benchmark for evaluating MLLMs’ geometric reasoning capabilities. Our datasets significantly and consistently improve the reasoning abilities of multiple MLLMs through both SFT and RL.
Future Work: We intend to extend NeSyGeo to other multimodal domains, such as analytical geometry and visual question answering. This extensibility will be achieved by defining new domain-specific languages, corresponding synthesis rules within the symbolic space, and tailored conversion engines. Furthermore, we plan to develop an automated symbolic solver capable of conducting search and validation directly within the symbolic space. This would remove reliance on LLMs, potentially reducing generation costs and ensuring complete correctness of the datasets.
Appendix A Comparison with Specific Examples of Popular Geometry Datasets
Figure 7: Comparison NeSyGeo-CoT dataset with other Popular Geometry Datasets. Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques. Our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
Appendix B Statistics of NeSyGeo-Caption and NeSyGeo-CoT

### v2

@@ 3.2 Symbolic Sequence Generation
Generator: To generate a Geo-DSL sequence, we employ a step-action augmentation procedure that iteratively synthesizes a sequence of statements, as detailed in Algorithm 1. Our methodology is initialized by configuring hyperparameters based on dataset characteristics: the sequence length $N$, weight matrices $\mathbf{I}$ and $\mathbf{A}$ that parameterize the selection distributions for geometric elements and symbolic actions (see Appendix H.2 for action details), and the value ranges $l_{\text{min}},l_{\text{max}}]$ and $\theta_{\text{min}},\theta_{\text{max}}]$ for lengths and angles, respectively.
The augmenter then commences an iterative synthesis over $N$ steps. At each step, after sampling continuous parameters ($x,y,z,\alpha$) and assigning randomized labels to the geometric vertices, our generative process selects a discrete element-action pair $(v_{j},a_{k})$ using a hierarchical weighted sampling strategy.
First, an element $v_{j}$ is sampled from the available set $f_{v}$. The probability is proportional to its weight from the matrix $\mathbf{I}$, modulated by an element selection temperature $\tau_{e}$:
$P(v_{j})=\frac{(\mathbf{w}_{I}(v_{j}))^{1/\tau_{e}}}{\sum_{j^{\prime}\in f_{v}}(\mathbf{w}_{I}(v_{j^{\prime}}))^{1/\tau_{e}}}$ (1)
Subsequently, conditioned on the chosen element $v_{j}$, a compatible action $a_{k}$ is sampled from the valid set $\mathcal{A}(v_{j})$, with its probability determined by weights from the matrix $\mathbf{A}$ and modulated by an action selection temperature $\tau_{a}$:
$P(a_{k}|v_{j})=\frac{(\mathbf{w}_{A}(a_{k},v_{j}))^{1/\tau_{a}}}{\sum_{k^{\prime}\in\mathcal{A}(v_{j})}(\mathbf{w}_{A}(a_{k^{\prime}},v_{j}))^{1/\tau_{a}}}$ (2)
The newly formed statement $s_{\text{new}}$ is then integrated into the sequence $f_{s}$. This two-stage, temperature-controlled probabilistic approach guarantees the validity of each statement while promoting significant diversity in the generated sequences.
3.3 CoT Generation
To generate Q&A pairs accompanied by reasoning paths, we first formalize the underlying deductive task. The objective is to navigate from the initial premises to a valid conclusion through a sequence of logical steps. This process can be conceptualized through the formulation of a reasoning graph:
$\mathcal{G}=(\mathcal{S},\mathcal{R},\hookrightarrow)$
where:
  • $\mathcal{S}$ is the set of all possible states, where each state $s\in\mathcal{S}$ represents the initial state or a deduced conclusion within the reasoning process.
  • $\mathcal{R}$ denotes the set of geometric rules, with each rule $r\in\mathcal{R}$ encoding a principle of logical deduction that governs the relationships between states.
  • $\hookrightarrow\subseteq\mathcal{S}\times\mathcal{R}\times\mathcal{S}$ is the state transition relation. A transition $(S_{r},r,s^{\prime})$ signifies that state $s^{\prime}$ is derived from a subset of states $S_{r}\subset\mathcal{S}$ by applying rule $r$.
However, a direct implementation of this formalism via a deterministic inference engine is fraught with significant limitations. Firstly, such an engine is computationally prohibitive, as it necessitates an exhaustive enumeration of the combinatorial space of conditions at each intermediate state. Given a state characterized by a set of $N$ known conditions, the engine must evaluate the applicability of $C$ fundamental axioms and theorems. For each rule $i$, which is predicated on matching a specific set of $c_{i}$ conditions, the search space grows combinatorially, with a complexity lower-bounded by $\mathcal{O}\left(\sum_{i=1}^{C}\binom{N}{c_{i}}\right)$, rendering any exhaustive search intractable for non-trivial geometric configurations. Secondly, this approach exhibits profound domain specificity. The reasoning framework and its constituent logical rules are intrinsically hard-coded for plane geometry. This tight coupling to a single domain severely curtails the model’s generalizability and fundamentally impedes its application for synthesizing datasets in other domains without substantial, non-trivial re-engineering.
To overcome these limitations, we pivot to leveraging the capabilities of LLMs, which can harness their extensive pre-trained knowledge and implicit intuitive faculties to facilitate rapid node expansion within the reasoning graph, obviating the need for bespoke rule engineering and thus offering a more agile and convenient paradigm. However, modern LLMs still face challenges when tackling intricate geometric tasks, and employing them for a direct, one-shot generation of Q&A pairs can result in erroneous information. To mitigate these issues, we introduce a novel two-stage methodology, which comprises a progressive reverse-search reasoner and a forward-validation verifier.
Reasoner. We introduce a novel search paradigm that operationalizes deductive reasoning as a Breadth-First Search (BFS) over the deductive lattice, accelerated by the learned preferences of LLMs. Instead of a uniform exploration, we leverage the reasoner to steer the expansion of the search frontier. From any given state $S_{t}$, the model $M_{\theta}$ is prompted to infer a single, contingent, and computable geometric conclusion $c_{t}$. This new conclusion augments the current state, effecting a state transition defined by $S_{t+1}=S_{t}\cup\{c_{t}\}$, where $c_{t}\leftarrow M_{\theta}(S_{t})$. This iterative expansion establishes a form of progressive supervision and effectively prunes the vast search space, ensuring the fidelity and diversity of the resultant reasoning trajectories. By decomposing the complex, long-horizon inference task into a sequence of incremental, verifiable deductions, it substantially lowers the reasoning complexity and mitigates the model’s propensity for hallucination. The terminal conclusions derived from this process are then randomly sampled to form candidate Q&A pairs.
Verifier. To verify the logical integrity of the generated Q&A pairs and to articulate a step-by-step CoT reasoning process. The questions generated by Reasoner, along with the corresponding conditions, are supplied as input. Verifier is prompted to produce both a final answer and its reasoning trajectory. These outputs are then cross-referenced against the answers derived from Reasoner. Only those pairs exhibiting concordance are retained for inclusion in our final dataset. The dual-verification protocol not only guarantees the correctness of the answers but also yields diverse and valid CoT narratives without necessitating an exhaustive traversal of the entire reasoning space.
@@ 3.4 Informalization
The formal symbolic sequences from our Geo-DSL must be mapped back into the visual and textual modalities. To compel cross-modal reasoning, we enforce a principle of information orthogonality: All explicit constraints and premises $I$ in initial state $S_{0}$are partitioned into disjoint visual modality $I_{v}$ and textual modality $I_{t}$ , such that $I_{v}\cap I_{t}=\emptyset$.
Painter. For the visual modality, painter parses each Geo-DSL statement as a geometric primitive, deterministically rendering high-fidelity images via Matplotlib. The rigorous symbolic space ensures the determinacy and uniqueness of the conversion process. Crucially, to mitigate inter-modal information imbalance, Painter employs a visual-centric annotation strategy, which involves embedding detailed information $I_{v}$ within the image space that are explicitly withheld from the accompanying textual description. Please refer to  A.2 for more details. This design choice robustly compels MLLMs to intrinsically leverage visual cues during problem-solving, thereby fostering enhanced visual perception and facilitating the extraction of perceptually-grounded information.
Translator. Translator converts each symbolic statement to multiple predefined natural language templates. To ensure diversity and randomness, each template is stochastically generated by GPT-4o and subsequently undergoes a manual curation process. This guarantees that every geometric element is associated with a minimum of eight corresponding descriptive templates. The final text condition avoids intermodal information redundancy, while diverse templates ensure the validity and diversity of the synthesized condition text.
@@ 4.1 Experimental Setup
Dataset. We generate a NeSyGeo-CoT dataset with 30k Q&A pairs and a NeSyGeo-Caption dataset with 70k Q&A pairs. The settings of generation process can be found in Appendix C.1. Additional dataset statistics and features are provided in Appendix B.1.
Evaluation. Our evaluation is conducted on several benchmarks: the Test set of GeoQA(Chen et al., 2021), the Test MINI set of MathVision(Wang et al., 2024b), and the MathVerse(Zhang et al., 2024c). For the MathVerse benchmark, we select the Vision Only, Vision Dominant, and Vision Intensive sets to better assess the visual perception and logical reasoning capabilities. We extract in-domain metrics from other datasets, including angle, area, length, and Plane Geometry, to effectively evaluate the models’ capabilities in geometric reasoning problems, in addition to the GeoQA dataset, which focuses entirely on plane geometry. For GeoQA, we employed hard-coded extraction for comparison, while other evaluations are assessed using the automated VLMEvalKit framework(Duan et al., 2024). Appendix C provides additional experimental details and evaluation results.
@@ 4.2 Empirical Results
Quality: Manually evaluated by human experts, our dataset was found to contain high-quality CoT rationales, accurate Q&A pairs, and aesthetically pleasing images.
Figure 4: Human evaluation results comparison.
To assess our dataset, we conducted a human expert evaluation with 40 experts (all at university undergraduate level or higher), comparing our NeSyGeo-CoT against G-LLaVA (Gao et al., 2025) and MAVIS(Zhang et al., 2024d). The evaluation was structured around five dimensions comprising nine metrics: Image, Question (including conditions), Reasoning path, Overall Impression and Final Answer. Each metric was rated on a 5-point scale. More details can be found in appendix C.2. NeSyGeo achieves state-of-the-art results on most metrics, ensuring a high degree of correctness, quality, and aesthetics, as corroborated by human evaluation.
To validate our two-stage generation strategy on CoT and Q&A quality, we also performed an ablation study on the Reasoner and Verifier. As shown in Table 1, we generated 50 instances for each configuration. The results demonstrate that the Reasoner successfully expands the deductive search space and reduces reasoning hallucinations, which ultimately enhances both the problem difficulty and the depth of the reasoning chains. Concurrently, the Verifier provides effective supervision that is instrumental in ensuring the logical soundness and high accuracy of the final data.
Table 1: Ablation study on the effects of the Reasoner and Verifier. Pass Rate measures the proportion of solutions accepted by the Verifier, while Pass@1 Rate (gpt-4o-mini’s) serves as a proxy for problem difficulty. All other metrics align with our human preference experiments.
Configurations Pass Rate CoT-Quality CoT-Accuracy Ans-Correctness Pass@1 Rate
Reasoner Verifier
X 78.0 4.34 4.66 5.00 68.0
X 4.40 4.22 4.42 52.0
64.0 4.68 4.54 5.00 48.0
Performance: Our dataset enhances the geometric reasoning capabilities of models through both RL or SFT. Under the same data budget, ours is better than that from popular automatic generation frameworks.
We first sample 4k samples from our NeSyGeo-CoT dataset and apply the Group Relative Policy Optimization (GRPO) algorithm to train two epochs with Deepseek R1’s format and answer rewards. As shown in Tables 2, InternVL2.5-4B significantly improved in the angle knowledge domain with gains of 8.4 (MathVerse), 7.3 (GeoQA), and 5.3 (MathVision). Qwen2.5-VL-3B achieved +15.8 performance boost in the area domain of MathVision. Notably, across all evaluated metrics, the InternVL2.5-4B model trained on the NeSyGeo dataset achieves performance on par with or superior to its 8B counterpart.
To ensure a fair comparison with others, we randomly sampled same budget from MAVIS (Zhang et al., 2024d) and R-CoT (Deng et al., 2024) and maintained consistent settings. As illustrated in Figure 5 and Table 2, while all datasets help the model improve over the baseline, our dataset outperformed others in most metrics, validating its superior performance.
@@
Model Angle Area Length Angle Area Length Plane
@@
GeoQA MathVerse
Model Angle Area Length Plane
@@
Qwen2.5-VL-7B+Ours 71.8 (+2.4) 46.1 (+3.1) 23.1 (-4.4) 49.5 (+3.3) 46.7 (+2.6)
LLaVA-7B 22.6 28.5 6.6 16.5 20.4
LLaVA-7B+Ours 26.1 (+3.5) 30.6 (+2.1) 7.7 (+1.1) 19.2 (+2.7) 22.9 (+2.5)
We also conducted SFT experiments, initially training on our NeSyGeo-Caption dataset to enhance the models’ perception of geometric images, followed by training on the NeSyGeo-CoT dataset to improve reasoning capabilities. Evaluation results on MathVerse (Vision Intensive) and GeoQA are presented in Table 3. The trained model demonstrates performance improvements over the base model on most metrics.
@@
Generability: The performance gains of models are from true reasoning, not just distilling solutions leaked from contaminated data
Although we have implemented stringent safeguards to ensure that our datasets do not access the test set, a critical inquiry is whether the efficacy of our framework stems from inadvertent data leakage introduced during the generation process. Adopting the methodology delineated by Wu et al. (Wu et al., 2025), we conducted an experiment utilizing the GeoQA test set.
For each question, we furnished the models with partial prompts, specifically truncated to the initial 40%, 60%, and 80% of their original length. We then evaluated two key metrics: (1) The ROUGE-L score, quantifying the textual overlap between the model’s generated output and gpt-4o’s output, and (2) The final answer accuracy. This empirical evaluation was performed on two distinct models, Qwen2.5-VL-3B and InternVL2.5-4B.
While a marginal increase in certain ROUGE-L scores as shown in Table 4, we posit that this is a manifestation of the model’s augmented reasoning capabilities, which empower it to perform more structured and relevant explorations from a limited context. The consistently negligible accuracy scores across all truncated inputs strongly indicate that the models have not merely memorized the test instance. Consequently, we conclude that there is no substantive evidence to suggest that the model’s performance gains are attributable to data contamination.
Table 4: Analysis of Potential Data Contamination using Truncated Prompts.
0.4 Ratio 0.6 Ratio 0.8 Ratio
Model ROUGE-L Acc. (%) ROUGE-L Acc. (%) ROUGE-L Acc. (%)
QwenVL-base 28.04 0.00 29.88 0.00 31.95 0.00
QwenVL-RL 21.29 0.00 36.72 0.00 40.57 0.00
InternVL-Base 26.82 0.00 40.04 0.00 48.07 4.69
InternVL-RL 30.00 0.00 44.91 0.00 57.37 1.54
Visual Perception: Models trained on NeSyGeo data shows modest gains over information-redundant datasets, yet achieves substantial improvements when image understanding is necessary. This indicates that ours enhance model’s cross-modal reasoning.
@@ 5 Conclusion
This paper introduces NeSyGeo, a neuro-symbolic framework for automatically synthesizing multimodal geometric datasets. Our approach transforms the generation process into a controllable symbolic space using Geo-DSL, utilizes Reasoner and Verifier for reverse search and forward validation to produce Q&A pairs, and then maps the symbolic representation back to image and natural language spaces via Painter and Translator. Using this framework, we construct NeSyGeo-CoT and NeSyGeo-Caption datasets, totalling 100k samples. We also propose NeSyGeo-Test, a comprehensive benchmark for evaluating MLLMs’ geometric reasoning capabilities. Our datasets significantly and consistently improve the reasoning abilities of multiple MLLMs through both SFT and RL.
Ethics Statement
NeSyGeo datasets presented in this paper are generated entirely through a synthetic, neuro-symbolic framework. The core generation process does not involve any user data, personally identifiable information, or web-scraping. Human involvement was limited exclusively to a voluntary study for evaluating the final dataset’s quality. All data collected from these human experts was fully anonymized to ensure strict confidentiality and the detailed evaluation protocol is provided in Appendix C.2. This methodology mitigates concerns related to both data privacy and social bias. Given the highly specialized nature of this mathematical task, we assess the potential for misuse or negative societal impact to be minimal. We release our contributions to the research community to responsibly foster progress in automated scientific reasoning.
Reproducibility Statement
To ensure the reproducibility of our research, we provide a comprehensive description of our data generation methodology throughout the main paper and in the Appendix A. Each module of our synthesis pipeline is delineated with its corresponding parameters and implementation details. We commit to making the complete synthetic dataset publicly available to the research community upon publication to facilitate further investigation and development. Furthermore, all hyperparameters and key configurations for our model training and experiments are fully disclosed in the Appendix C.
Appendix Contents
Appendix A More Details of Methods
A.1 Generator
As outlined in Algorithm 1, for each step, we first generate three lengths, $x$, $y$, and $z$, along with an angle $\alpha$, sampled from predefined ranges. We then randomly selecting vertex labels. Subsequently, based on the weight matrices $A$ and $I$, we determine the specific statement to be selected. The chosen statement, paired with its corresponding numerical values, is then appended to the Geo-DSL sequence.
1:Input: Step count NN, Action weight matrix $A$, Element selection weight matrix $I$, Line length range $l_{min},l_{max}]$, Angle range $\theta_{min},\theta_{max}]$$\triangleright$ Set customizable hyperparameter
2:Output: Generated Geo-DSL statement sequence $f_{s}$.
5:for $i=1$ to NN do
6:  Randomly sample $x,y,z$ from $l_{min},l_{max}]$.
7:  Randomly sample $\alpha$ from $\theta_{min},\theta_{max}]$.
8:  Element $v_{j}$=Selected_elements($f_{v}$,II). $\triangleright$ Select element $v_{j}$ from $f_{v}$ randomly.
9:  Action $a_{k}$=Selected_action($v_{j}$,AA). $\triangleright$ Select action $a_{k}$ based on the type of $v_{j}$ randomly
11:  $f_{s}$=Update($f_{s}$$s_{new}$) $\triangleright$ Add the new statement to the sequence
12:  $f_{v}$=Update($f_{v}$$s_{new}$). $\triangleright$ Update state space elements
13:end for
14:return $f_{s}$.
A.2 Painter
A primary challenge in machine-based geometric analysis is the frequent ambiguity between visual diagrams and the text that describes them. This disconnect between the visual and textual data streams creates a significant impediment for models attempting to perform integrated, cross-modal reasoning. To address this, we developed Painter, a rendering protocol that systematically embeds explicit semantic information directly into the geometric figures themselves.
Painter ensures that each visual element has a clear and unambiguous interpretation by applying a strict set of annotation rules. The strategies are as follows:
  • Segment Length: Each line segment is explicitly labeled with its identifier and corresponding numerical length, with the value rounded to one decimal place.
  • Angle Annotations: The measure of any non-right angle is directly inscribed within the angle’s arc. For right angles, the conventional square symbol is used exclusively.
  • Circle Radius: The radius of a circle is represented by a dashed line from the center to a point on the circumference, clearly marked with the label “r=[length]’‘.
To further enhance the visual diversity and aesthetic quality of the generated figures, Painter incorporates several stochastic rendering strategies. For each synthesis, the global scale ( the pixel-to-unit ratio) is randomly sampled, and the geometric figure is randomly translated within the canvas. Additionally, distinct classes of geometric elements are rendered using a varied color palette to improve their visual discriminability. These procedures ensure a high degree of variation in the training data, preventing the model from overfitting to superficial stylistic features such as position or scale.
Crucially, to prevent information redundancy and compel a true synthesis of visual and textual information, any property that is graphically rendered by Painter is intentionally omitted from the accompanying text description.
A.3 Translator
To provide the Reasoner with sufficient context for its reasoning process, our Translator module initially generates a complete and comprehensive set of textual conditions. However, to mitigate information inconsistency across modalities in the final synthesized dataset, the Translator performs a crucial refinement step. It systematically prunes any explicit constraint-based information that is visually rendered in the image (e.g., numerical values for segment lengths or angles). This process prevents the model from developing “shortcuts” based on redundant textual data, which would impair its cross-modal reasoning capabilities. The final text exclusively retains abstract information about entities (e.g., ”isosceles triangle”) and their semantic relations (e.g., ”point O is the incenter of triangle ABC”), compelling the model to ground its understanding in the visual domain.
A.4 Reasoner
We employ DeepSeek-R1 as our reasoning LLM to conduct a forward search process. At each step, the model receives the conditions of the current state and is prompted to infer a single, new contingent and computable geometric property or relationship that augments this state. The specific prompt used for this task is shown in Box A.4.
The prompt for Reasoner to reverse search Use the mathematics you know to make simple and accurate inferences and get the conclusions based on image descriptions. Focus on extracting as much information as possible. Keep the final answer of all reasoning to one decimal place. Use <<conclusion>> <</conclusion>> to include all your inferences. Example is as follows: Output: <<conclusion>> 1. The area of Triangle SFD is 18.0. 2. The length of line EF is 3.9. 3. The measure of angle DFS is 29.4°. <</conclusion>> Here is the image description input: Current state
A.5 Verifier
We employ DeepSeek-V3 to conduct the forward verification process. The prompt used for this task is detailed in Box A.5.
The prompt for the Verifier to Forward Validation Here is a geometry problem with the following information: Image description: Initial state Question: Questions from Reasoner Following the above condition and question, think step by step and answer the following question directly. After thinking process, you should provide your final concise reasoning steps in <<steps>><</steps>> tags and your final answer in <<answer>> </</answer>> tags. The answer to the question is only allowed to contain one number or angle.
Appendix B More Details of Datasets
B.1 Statistics of Datasets